OpenStudy (anonymous):

Solve. p + 11 – 4p = 8 p= ?

Mehek (mehek14):

ok first add p + (-4p)

OpenStudy (anonymous):

I don't think I understand. I feel kind of stupid. Sorry.

Mehek (mehek14):

p + (-4p) = -3p

OpenStudy (anonymous):

Okay.

OpenStudy (dmndlife24):

add like terms p -4p = -3p now subtract 11 on both sides to to get the p terms by itself... -3p = -3 p = ?

OpenStudy (anonymous):

-14?

OpenStudy (dmndlife24):

What is -3 divided by -3 ?????????

OpenStudy (anonymous):

1

OpenStudy (dmndlife24):

Yes, we check by plugging 1 in for p... (1) + 11 - 4(1) = 8 8 = 8 :D

OpenStudy (anonymous):

Thank you. :) Sorry for being stupid.

OpenStudy (dmndlife24):

Lol, your welcome